### v1

@@ Abstract
Resources for simulation-based evaluation of conversational recommender systems (CRSs) are scarce. The UserSimCRS toolkit was introduced to address this gap. In this work, we present UserSimCRS v2, a significant upgrade aligning the toolkit with state-of-the-art research. Key extensions include an enhanced agenda-based user simulator, introduction of large language model-based simulators, integration for a wider range of CRSs and datasets, and new LLM-as-a-judge evaluation utilities. We demonstrate these extensions in a case study.
@@ 1 Introduction
There is a growing interest in conversational recommender systems (CRSs), which are designed to support users in finding items corresponding to their needs and preferences through a multi-turn dialogue [13]. Despite the progress in the field, CRS evaluation remains a challenging task [13, 17, 4]. Indeed, traditional evaluation methods using offline test collections do not capture the interactive nature of the systems. To address this issue, evaluation methods relying on humans are often employed, yet they suffer from significant limitations, such as their reproducibility, scalability, and cost [18]. Therefore, user simulation has recently been gaining interest to mitigate these limitations [38, 4, 7, 26, 35, 16, 9, 44, 32, 41, 5]. While user simulation can be imperfect, it serves as a valuable preliminary step before human evaluation to refine and make a pre-selection of the best-performing systems, thus optimizing the use of human resources [1].
@@
  • Advanced conversational quality evaluation: We expand the toolkit’s limited evaluation metrics by adding a new “LLM-as-a-Judge” utility.
@@ 3 UserSimCRS v1
While the toolkit is a valuable asset for the community, it has some limitations that may hinder its widespread adoption.
@@
  • Lack of support for benchmark datasets: The toolkit provides a dataset to create a user simulator, but it is not widely used in the field. Hence, the preparation of popular benchmark datasets is left to the toolkit users which represents a significant burden.
@@ 4 UserSimCRS v2 Extensions
The underlying objective of UserSimCRS v2 remains the same as its predecessor: to facilitate the comparison of different conversational recommender systems via simulation-based evaluation. Additionally, it can be used to compare and investigate different user simulators. In this section, we present the extensions introduced in UserSimCRS v2, which are illustrated in Fig. 2. Specifically, we describe the integration of existing benchmark datasets (Section 4.1) the improvements to the agenda-based user simulator (Section 4.2), the introduction of LLM-based user simulators (Section 4.3), the integration of existing CRSs (Section 4.5), and the addition of an evaluation utility to assess the conversation quality (Section 4.6).
@@ 4.1 Unified Data Format and Benchmark Datasets
The proposed unified data format is based on the format used in DialogueKit, where utterances are annotated with dialogue acts. A dialogue act comprises an intent and optionally a set of associated slot-value pairs. For example, the dialogue acts Elicit(size) and Elicit(color) are present in the following CRS utterance: “What size and color do you prefer for your new skates?” This allows for a refined representation of utterances and to support more complex interactions than UserSimCRS v1 which only considers a single intent and a list of slot-value pairs per utterance. Furthermore, we introduce the notion of an information need that consists of a set of constraints $C$ and requests $R$ as in [27], in addition to target items (i.e., what the user is looking for). For example, the information need of a user looking for the song Happy by Pharrell Williams and who wants to know the release year can be represented as follows:
@@
We note that these benchmark datasets do not necessarily contain all the annotations required to build and train certain user simulators. For example, the utterances in ReDial are not annotated with dialogue acts which are needed to train a user simulator or some components of it (e.g., a supervised natural language understanding model). In such cases, we propose to perform automatic data augmentation. Previous work [28, 40] argues that LLMs are an acceptable alternative to human annotators in data-scarce scenarios. Therefore, in UserSimCRS v2, we provide tools relying on pre-trained LLMs for the tasks of dialogue act annotation and information need extraction. However, when using these tools, the biases and limitations of LLMs should be acknowledged.
@@ 4.2 Enhanced Agenda-based User Simulator
The objective of the NLU component is to extract dialogue acts from incoming CRS utterances. In UserSimCRS v2, we implement a LLM-based NLU component which is initialized with a list of possible intents and slots for the dialogue acts and a prompt that includes a placeholder for the incoming CRS utterance. Dialogue acts extraction is performed by sending the prompt with a given utterance to an LLM (hosted on Ollama [^3] https://ollama.com) and parsing the output based on the following expected format: intent1(slot="value",slot, ...)|intent2(). We provide a default prompt, built using a few-shot approach with examples taken from the augmented IARD dataset, to extract dialogue acts in the movie domain, along with its corresponding configuration.
@@
The NLG component is responsible for generating human-like textual utterance based on the dialogue acts provided by the dialogue policy component. In UserSimCRS v2, we introduce an LLM-based NLG component (relying on an LLM hosted on Ollama). It aims to increase the diversity of the generated responses compared to the template-based NLG component from UserSimCRS v1. To generate an utterance, this component uses a prompt that includes placeholders for dialogue acts and, optionally, additional annotations (e.g., emotion). We also provide a default prompt, following a few-shot approach with examples from the augmented IARD dataset, for the movie domain, alongside its corresponding configuration.
@@ 4.4 Integration with Large Language Models
UserSimCRS v2 introduces several components and user simulators relying on an LLM, all of which assumes the LLM employed is hosted on a server. To manage interactions with the server, all LLM-based modules use a common interface. UserSimCRS v2 provides two initial implementations of this interface: one for OpenAI [^4] https://openai.com/api/ and one for an Ollama server. This design makes the toolkit highly extensible, as additional interfaces for other LLM backends can be easily added through inheritance.
@@ 4.5 Integration with Existing CRSs
4.6 LLM-as-a-Judge for Conversation Quality Evaluation
Evaluation in UserSimCRS v1 is limited to user satisfaction and average number of turns. However, there are additional measures to assess the performance of CRSs, such as those related to recommendation accuracy and linguistic abilities [17]. In UserSimCRS v2, we provide a new evaluation utility to assess conversation quality using a LLM-based evaluator (“LLM-as-a-Judge”). Conversation quality is assessed with respect to five aspects inspired by Huang et al. [16]: (1) Recommendation relevance measures how closely the recommended items align with the user’s preferences and needs; (2) Communication style corresponds to the conciseness and clarity of the responses; (3) Fluency is the degree of naturalness of the responses compared to human-generated responses; (4) Conversational flow assesses the coherence and consistency of the conversation; and (5) Overall satisfaction encapsulates the user’s holistic experience. Each aspect is assigned a score between 1 and 5, with each score defined in a grading rubric. [^5] The grading rubrics are available at: https://github.com/iai-group/UserSimCRS/blob/main/scripts/evaluation/rubrics/quality˙rubrics.py For example, for recommendation relevance a score of 1 represents irrelevant recommendations and a score of 5 represents highly relevant recommendations. We acknowledge that there are open questions regarding the use of LLM-based evaluators, as correlation with human judgments varies across different studies [11].
@@ 6 Conclusion
In this work, we presented an upgraded version of UserSimCRS, designed to facilitate a more comprehensive and flexible comparison of user simulators and conversational recommender systems through simulation-based evaluation. This new version aligns the toolkit with current research by enhancing the classic agenda-based user simulator with LLM-powered components, introducing LLM-based simulators, adding new LLM-based automatic evaluation utilities, and providing support for a wider range of CRSs and datasets.

### v2

@@ Abstract
Resources for simulation-based evaluation of conversational recommender systems (CRSs) are scarce. The UserSimCRS toolkit was introduced to address this gap. In this work, we present UserSimCRS v2, a significant upgrade aligning the toolkit with state-of-the-art research. Key extensions include an enhanced agenda-based user simulator, introduction of large language model-based simulators, integration for a wider range of CRSs and datasets, and new evaluation utilities. We demonstrate these extensions in a case study.
@@ 1 Introduction
There is a growing interest in conversational recommender systems (CRSs), which are designed to support users in finding items corresponding to their needs and preferences through a multi-turn dialogue [13]. Despite the progress in the field, CRS evaluation remains a challenging task [13, 17, 4, 5]. Indeed, traditional evaluation methods using offline test collections do not capture the interactive nature of the systems. To address this issue, evaluation methods relying on humans are often employed, yet they suffer from significant limitations, such as their reproducibility, scalability, and cost [18]. Therefore, user simulation has recently been gaining interest to mitigate these limitations [38, 4, 7, 26, 35, 16, 9, 44, 32, 41, 5]. While user simulation can be imperfect, it serves as a valuable preliminary step before human evaluation to refine and make a pre-selection of the best-performing systems, thereby optimizing the use of human resources [1].
@@
  • Advanced conversational quality evaluation: We expand the toolkit’s limited evaluation metrics by adding a new “LLM-as-a-Judge” utility along with the implementation of recently proposed user-centric utility metrics [5].
@@ 3 UserSimCRS v1
While the toolkit is a valuable asset for the community, we believe that the following limitations hinder its widespread adoption.
@@
  • Lack of support for benchmark datasets: The toolkit provides a dataset based on the IAI MovieBot agent [14] to create a user simulator, but it is not widely used in the field. Hence, the preparation of popular benchmark datasets is left to the toolkit users which represents a significant burden.
@@ 4 UserSimCRS v2 Extensions
The underlying objective of UserSimCRS v2 remains the same as its predecessor: to facilitate the comparison of different conversational recommender systems via simulation-based evaluation. Additionally, it can be used to compare and investigate different user simulators. In this section, we present the extensions introduced in UserSimCRS v2, which are illustrated in Fig. 2. Specifically, we describe the integration of existing benchmark datasets (Section 4.1) the improvements to the agenda-based user simulator (Section 4.2), the introduction of LLM-based user simulators (Section 4.3), the integration of existing CRSs (Section 4.5), and the addition of evaluation utilities and metrics to assess conversation quality and user utility (Section 4.6).
@@ 4.1 Unified Data Format and Benchmark Datasets
The proposed unified data format is based on the format used in DialogueKit, where utterances are annotated with dialogue acts. A dialogue act comprises an intent and optionally a set of associated slot-value pairs. For example, the dialogue acts Elicit(size) and Elicit(color) are present in the following CRS utterance: “What size and color do you prefer for your new skates?” This allows for a refined representation of utterances and to support more complex interactions than UserSimCRS v1, which only considers a single intent and a list of slot-value pairs per utterance. Furthermore, we introduce the notion of an information need that consists of a set of constraints $C$ and requests $R$ as in [27], in addition to target items (i.e., what the user is looking for). For example, the information need of a user looking for the song Happy by Pharrell Williams and who wants to know the release year can be represented as follows:
@@
We note that these benchmark datasets do not necessarily contain all the annotations required to build and train certain user simulators. For example, the utterances in ReDial are not annotated with dialogue acts which are needed to train a user simulator or some components of it (e.g., a supervised natural language understanding model). In such cases, we propose to perform automatic data augmentation. Previous work [28, 40] argues that LLMs are an acceptable alternative to human annotators in data-scarce scenarios. Therefore, in UserSimCRS v2, we provide tools relying on pre-trained LLMs for the tasks of dialogue act annotation and information need extraction. However, when using these tools, the biases and limitations of LLMs should be acknowledged and sanity checks are recommended.
@@ 4.2 Enhanced Agenda-based User Simulator
The objective of the NLU component is to extract dialogue acts from incoming CRS utterances. In UserSimCRS v2, we implement a LLM-based NLU component which is initialized with a list of possible intents and slots for the dialogue acts and a prompt that includes a placeholder for the incoming CRS utterance. Dialogue acts extraction is performed by sending the prompt with a given utterance to an LLM and parsing the output based on the following expected format: intent1(slot="value",slot, ...)|intent2(). We provide a default prompt, built using a few-shot approach with examples taken from the augmented IARD dataset, to extract dialogue acts in the movie domain, along with its corresponding configuration.
@@
The NLG component is responsible for generating human-like textual utterance based on the dialogue acts provided by the dialogue policy component. In UserSimCRS v2, we introduce an LLM-based NLG component. It aims to increase the diversity of the generated responses compared to the template-based NLG component from UserSimCRS v1. To generate an utterance, this component uses a prompt that includes placeholders for dialogue acts and, optionally, additional annotations (e.g., emotion). We also provide a default prompt, following a few-shot approach with examples from the augmented IARD dataset, for the movie domain, alongside its corresponding configuration.
@@ 4.4 Integration with Large Language Models
UserSimCRS v2 introduces several configurable components and user simulators relying on an LLM, all of which assumes the LLM employed is hosted on a server. To manage interactions with the server, all LLM-based modules use a common interface. UserSimCRS v2 provides two initial implementations of this interface: one for OpenAI [^3] https://openai.com/api/ and one for an Ollama [^4] https://ollama.com/ server. This design makes the toolkit highly extensible, as additional interfaces for other LLM backends can be easily added through inheritance.
@@ 4.5 Integration with Existing CRSs
4.6 Advanced Evaluation Metrics
Evaluation in UserSimCRS v1 is limited to user satisfaction and average number of turns. However, there are additional measures to assess the performance of CRSs, such as those related to recommendation accuracy and linguistic abilities [17]. In UserSimCRS v2, we provide new evaluation utilities to assess conversation quality using a LLM-based evaluator (“LLM-as-a-Judge”). We also implement novel metrics that measure utility from a user-centric perspective [5].
Conversation Quality Evaluation with LLM-as-a-Judge.
Conversation quality is assessed with respect to five aspects inspired by Huang et al. [16]: (1) Recommendation relevance measures how closely the recommended items align with the user’s preferences and needs; (2) Communication style corresponds to the conciseness and clarity of the responses; (3) Fluency is the degree of naturalness of the responses compared to human-generated responses; (4) Conversational flow assesses the coherence and consistency of the conversation; and (5) Overall satisfaction encapsulates the user’s holistic experience. Each aspect is assigned a score between 1 and 5, with each score defined in a grading rubric. [^5] The grading rubrics are available at: https://github.com/iai-group/UserSimCRS/blob/main/scripts/evaluation/rubrics/quality˙rubrics.py For example, for recommendation relevance a score of 1 represents irrelevant recommendations and a score of 5 represents highly relevant recommendations. We acknowledge that there are open questions regarding the use of LLM-based evaluators, as correlation with human judgments varies across different studies [11].
User-centric Utility Metrics
. Bernard and Balog [5] propose metrics to assess the utility of CRSs from a user-centric perspective. We implement Successful Recommendation Round Ratio (SRRR) and Reward-per-Dialogue-Length (RDL), in addition to success rate. SRRR measures the proportion of recommendation rounds resulting in an acceptance, while RDL divides the number of accepted recommendations by dialogue length. The computation of these metrics relies on the underlying intent of user utterances in a conversation, that are annotated either manually or automatically using an NLU model.
@@ 6 Conclusion
In this work, we presented an upgraded version of UserSimCRS, designed to facilitate a more comprehensive and flexible comparison of user simulators and conversational recommender systems through simulation-based evaluation. This new version aligns the toolkit with current research by enhancing the classic agenda-based user simulator with LLM-powered components, introducing LLM-based simulators, adding new automatic evaluation utilities, and providing support for a wider range of CRSs and datasets.
@@ 6.0.1 Acknowledgements
6.0.2 Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.